# Schatten classes on noncommutative tori: Kernel conditions

### v1

@@ 2.1. Noncommutative Tori
Harmonic analysis on noncommutative tori is an established subject. The exposition here closely follows [13], and for the sake of brevity, we refer the reader to [13] for a detailed exposition of the topic and provide here only the definitions relevant to this text.
@@ 2.1.2. Harmonic analysis on quantum tori
2.2. Schatten classes and Schur multipliers
@@ 3.1. Schwartz’ kernel theorem
$\langle Tx,y\rangle=\langle T_{k}x,y\rangle=\langle k,y\otimes x\rangle.$
@@
(3.2) $A(x,y)=\langle\omega,x\otimes y\rangle$
@@
$\langle\omega,h\rangle=\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n}),$
@@
$\begin{split}|\langle\omega,h\rangle|&=|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|\\ &\leq\sum_{m,n\in\mathbb{Z}^{d}}|a_{m,n}||A(V^{m},V^{n})|\\ &\lesssim\sum_{m,n\in\mathbb{Z}^{d}}\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}}(1+|m|^{2})^{-\frac{s_{0}}{2}}(1+|n|^{2})^{-\frac{s_{0}}{2}}\|A\|\\ &\leq C\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}},\end{split}$
@@
(3.7) $\begin{split}\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)&={\rm{sup}}\{|A(x,y)|,\ x,y\in B_{{\mathbb{T}}_{\theta}^{d}}\},\\ \rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)&={\rm{sup}}\{|\langle\omega,h\rangle|,\,h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}\},\end{split}$
@@
(3.8) $\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(\Lambda\omega)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|A(x,y)|=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|\langle\omega,x\otimes y\rangle|.$
@@
$\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|\langle\omega,x\otimes y\rangle|\leq\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\langle\omega,h\rangle|=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega),$
@@
$\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\Lambda^{-1}A)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\langle\omega,h\rangle|=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|.$
@@
$\langle Tx,y\rangle=\langle T_{k}x,y\rangle\ \text{for}\ x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d}),$
where $T_{k}$ is the operator associated to the kernel $k$ implemented by $\langle T_{k}x,y\rangle=\langle k,y\otimes x\rangle$ for $x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d})$.
Now for any $\theta$ and $1\leq p\leq\infty$, the space $L^{p}({\mathbb{T}}_{\theta}^{d})$ embeds into $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$: an element $x\in L^{p}({\mathbb{T}}_{\theta}^{d})$ derives a continuous functional on $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ by $y\mapsto\tau(xy)$. A map $T\in B(L^{2}({\mathbb{T}}_{\theta}^{d}))$ can be viewed as a linear map from $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ to $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$, by plugging into Theorem 3.1, we obtain that there exists $k\in\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ such that
$\langle Tx,y\rangle=\langle k,y\otimes x\rangle.$
@@
(3.9) $\begin{split}\langle Tx,y\rangle&=\langle T_{k}x,y\rangle\\ &=\langle k,y\otimes x\rangle\\ &=\tau\otimes\tau(k(y\otimes x))\\ &=\tau(({\rm id}\otimes\tau)(k(1\otimes x))y),\end{split}$

### v2

@@ 2.1. Noncommutative Tori
Harmonic analysis on noncommutative tori is an established subject. The exposition here closely follows [6] and [13], and for the sake of brevity, we refer the reader to [6] and [13] for a detailed exposition of the topic and provide here only the definitions relevant to this text.
@@ 2.1.1. Basic definitions
There is an alternative way to understand the noncommutative tori, given $m,n\in\mathbb{Z}^{d}$, define a map from $\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ to ${\mathbb{T}}$:
$\sigma(m,n)=e^{2\pi{\rm{i}}m^{t}\,\tilde{\theta}\,n}.$
Here $m^{t}$ denotes the transpose of $m=(m_{1},\dots,m_{d})$ and $\tilde{\theta}$ is the following lower-triangular $d\times d$-matrix deduced from the skew symmetric matrix $\theta$:
(2.2) $\tilde{\theta}=\begin{pmatrix}0&0&0&\dots&0\\ \theta_{2,1}&0&0&\dots&0\\ \theta_{3,1}&\theta_{3,2}&0&\dots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \theta_{d-1,1}&\theta_{d-1,2}&\theta_{d-1,3}&\dots&0\\ \theta_{d,1}&\theta_{d,2}&\theta_{d,3}&\dots&0\end{pmatrix}.$
Define a projective representation of $\mathbb{Z}^{d}$ on $\ell_{2}(\mathbb{Z}^{d})$ induced by the above $\sigma$:
$\lambda_{\sigma}(m)g(n)=\sigma(-n,m)g(m-n).$
Then, for $m,n\in{\mathbb{Z}}^{d}$, we have
(2.3) $\lambda_{\sigma}(m)\lambda_{\sigma}(n)=\sigma(m,n)\lambda_{\sigma}(m+n).$
Consider the twisted group von Neumann algebra $\mathcal{L}_{\sigma}({\mathbb{Z}}^{d})$ generated by $\{\lambda_{\sigma}(m),\ m\in{\mathbb{Z}}^{d}\}$, and a direct computation shows that
$U^{m}U^{n}=\sigma(m,n)U^{m+n}.$
By Stone-von Neumann theorem, $\mathcal{L}_{\sigma}({\mathbb{Z}}^{d})$ is *-isomorphic to ${\mathbb{T}}_{\theta}^{d}$ by mapping $U^{m}$ to $\lambda_{\sigma}(m)$ for $m\in{\mathbb{Z}}^{d}$.
The tracial state $\tau_{\sigma}$ on $\mathcal{L}_{\sigma}({\mathbb{Z}}^{d})$ is induced by the cyclic vector $\delta_{0}$, which is the dirac function taking value at the unital element $(0,...,0)$ of ${\mathbb{Z}}^{d}$. Namely, for any $x\in\mathcal{L}_{\sigma}({\mathbb{Z}}^{d})$, we obtain that
$\tau_{\sigma}(x)=\langle x\delta_{0},\delta_{0}\rangle.$
One can verify that for $x=\sum_{m\in{\mathbb{Z}}^{d}}\alpha_{m}\lambda_{\sigma}(m)$ with finite sum, we have $\tau_{\sigma}(x)=\alpha_{0}$. Which indicates that the two traces $\tau$ and $\tau_{\sigma}$ coincide.
Moreover, by the structure of the twisted group algebra over ${\mathbb{Z}}^{d}$. We can compute the matricial representation of an element $x\in{\mathbb{T}}_{\theta}^{d}$. Namely, with the twisted convolution and involution on $\ell_{1}({\mathbb{Z}}^{d})$:
(2.4) $f*_{\sigma}g(m)=\sum_{n\in{\mathbb{Z}}^{d}}f(m-n)g(n)\sigma(m-n,n),$
(2.5) $f^{\sharp}(m)=\sigma(m,-m)^{*}\overline{f(-m)}.$
We gain that
$\lambda_{\sigma}(f)\lambda_{\sigma}(g)=\lambda_{\sigma}(f*_{\sigma}g)$
and
$\lambda_{\sigma}(f)g=f*_{\sigma}g,$
for $f,g\in\ell_{1}({\mathbb{Z}}^{d})$.
Now, we turn to another important perspective for computing the representation of $x\in{\mathbb{T}}_{\theta}^{d}$, the noncommutative Fourier transform.
@@
We will consider the left multiplication operators $M_{x}$ by mapping $y\in L_{2}({\mathbb{T}}_{\theta}^{d})$ to $xy$ as operators on the Hilbert space $\ell_{2}(\mathbb{Z}^{d})$ by virtue of the Plancherel formula (2.6). Take a left multiplication $M_{x}$, where $x\sim\sum_{m\in\mathbb{Z}^{d}}\widehat{x}(m)U^{m}$. By (2.6), any $\eta\in L_{2}({\mathbb{T}}^{d}_{\theta})$ corresponds to the element $\{\widehat{\eta}(n)\}_{n\in{\mathbb{Z}}^{d}}\in\ell_{2}({\mathbb{Z}}^{d})$. By (2.5) and (2.4), a direct computation shows that the Fourier coefficient of $M_{x}\eta$ at $m\in{\mathbb{Z}}^{d}$ is
$\sum_{n\in{\mathbb{Z}}^{d}}\sigma(m-n,n)\widehat{x}(m-n)\widehat{\eta}(n).$
Therefore, we can represent the operator $M_{x}$ as a matrix indexed by ${\mathbb{Z}}^{d}$:
(2.7) $\Big{(}\sigma(m-n,n)\widehat{x}(m-n)\Big{)}_{m,n\in{\mathbb{Z}}^{d}}.$
@@ 2.1.2. Harmonic analysis on quantum tori
In the same spirit with (2.7), Fourier multipliers on ${\mathbb{T}}^{d}_{\theta}$ may be regarded as diagonal matrices acting on $\ell_{2}({\mathbb{Z}}^{d})$. For instance, the potentials $I^{\alpha}$ and $J^{\alpha}$ may be represented as
(2.9) ${\rm{diag}}\{|m|^{\alpha}\}_{m\in{\mathbb{Z}}^{d}}\quad\mbox{ and }\quad{\rm{diag}}\{(1+|m|^{2})^{\frac{\alpha}{2}}\}_{m\in{\mathbb{Z}}^{d}}$
respectively.
@@
2.2. Schatten classes
@@ 3.1. Schwartz’ kernel theorem
Before delving into the main parts of this note, we clarify some notations.
[…]
@@
$(Tx,y)=(T_{k}x,y)=(k,y\otimes x).$
@@
(3.2) $A(x,y)=(\omega,x\otimes y)$
@@
$(\omega,h)=\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n}),$
@@
$\begin{split}|(\omega,h)|&=|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|\\ &\leq\sum_{m,n\in\mathbb{Z}^{d}}|a_{m,n}||A(V^{m},V^{n})|\\ &\lesssim\sum_{m,n\in\mathbb{Z}^{d}}\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}}(1+|m|^{2})^{-\frac{s_{0}}{2}}(1+|n|^{2})^{-\frac{s_{0}}{2}}\|A\|\\ &\leq C\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}},\end{split}$
@@
(3.7) $\begin{split}\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)&={\rm{sup}}\{|A(x,y)|,\ x,y\in B_{{\mathbb{T}}_{\theta}^{d}}\},\\ \rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)&={\rm{sup}}\{|(\omega,h)|,\,h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}\},\end{split}$
@@
(3.8) $\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(\Lambda\omega)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|A(x,y)|=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|(\omega,x\otimes y)|.$
@@
$\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|(\omega,x\otimes y)|\leq\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|(\omega,h)|=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega),$
@@
$\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\Lambda^{-1}A)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|(\omega,h)|=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|.$
@@
$(Tx,y)=(T_{k}x,y)\ \text{for}\ x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d}),$
where $T_{k}$ is the operator associated to the kernel $k$ implemented by $(T_{k}x,y)=(k,y\otimes x)$ for $x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d})$.
Now for any $\theta$ and $1\leq p\leq\infty$, the space $L^{p}({\mathbb{T}}_{\theta}^{d})$ embeds into $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$: an element $x\in L^{p}({\mathbb{T}}_{\theta}^{d})$ derives a continuous functional on $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ by taking $(x,y)=\tau(xy)$. A map $T\in B(L^{2}({\mathbb{T}}_{\theta}^{d}))$ can be viewed as a linear map from $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ to $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$, by plugging into Theorem 3.1, we obtain that there exists $k\in\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ such that
$(Tx,y)=(k,y\otimes x).$
@@
(3.9) $\begin{split}(Tx,y)&=(T_{k}x,y)\\ &=(k,y\otimes x)\\ &=\tau\otimes\tau(k(y\otimes x))\\ &=\tau(({\rm id}\otimes\tau)(k(1\otimes x))y),\end{split}$